OpenStudy (anonymous):

is the line through the points R(-1,3) and S(2,-7) parallel to the graph of the line given by the equation, 10x + 3y=6

OpenStudy (anonymous):

A) no, the slopes have opposite signs B) yes, they have the same sign and value C) Yes, the lines both decrease to the right D) No, the lines have unequal slopes

OpenStudy (luigi0210):

Start by finding the slope

OpenStudy (anonymous):

\[slope of RS=\frac{-7-3 }{2+1 }=\frac{ -10 }{3}\]

OpenStudy (luigi0210):

Well that's done now xD Okay, so now use the point-slope formula.

OpenStudy (anonymous):

i don't kow what that is..

OpenStudy (luigi0210):

Justina! o: \[\huge y-y_{1}=m(x-x_{1})\]

OpenStudy (anonymous):

I sorry Jesus!! I'm really bad at this

OpenStudy (luigi0210):

It's alright, do you know what the equations means?

OpenStudy (anonymous):

no not really..

OpenStudy (luigi0210):

The point-slope formula corresponds with the points: \[\LARGE (x_{1}, y_{1})\]

OpenStudy (anonymous):

okay

OpenStudy (luigi0210):

and m=slope, so just plug everything in: \[\LARGE y-3=-\frac{10}{3}(x+1)\]

OpenStudy (anonymous):

okay so how do i solve this?

OpenStudy (luigi0210):

Distribute :)

OpenStudy (luigi0210):

Did you ever learn if it was parallel or not? >.>

OpenStudy (anonymous):

No i'm trying to figure it out right now haha

OpenStudy (luigi0210):

You know what Justina, slap me right now -.- I made this problem harder than it needed to be ._.

OpenStudy (anonymous):

haha i would never slap yoU!

OpenStudy (luigi0210):

but yea, we should of stopped after we found slope ._.

OpenStudy (anonymous):

Okay so i'm confused now

OpenStudy (luigi0210):

Okay, so when finding parallel lines, all we need to do is determine whether the SLOPES are the same. So the slope for the points was : \[\LARGE m=-\frac{10}{3}\] After we found that we should of solved the second equation: \[\LARGE 10x+3y=6\] \[\LARGE 3y=-10x+6\] \[\LARGE y=-\frac{10}{3}x+2\] And both have the -10/3 So they are parallel .-.

OpenStudy (anonymous):

they are parallel because the slopes have the same sign and value?

OpenStudy (luigi0210):

Yup

OpenStudy (luigi0210):

Sorry about that Justina .-.

OpenStudy (anonymous):

Ohh okay!! it's okay Jesus! it makes sense now!

OpenStudy (luigi0210):

Alright, if you say so >.> I'm an awful teacher D:

OpenStudy (anonymous):

no it's okay!! no worries you still helped me out!